# Collaborative and parametric insurance on the Ethereum blockchain

### v1

@@ 2.2.1 Parametric insurance portfolio over time
A parametric insurance portfolio is a collection of parametric insurance contracts $\{(A_{T_{i}},l_{i})\}_{i\geq 1}$ defined on the probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$. Let $S_{i}$ denote the time at which contract $i$ is underwritten. The number of contracts underwritten up to time $t\geq 0$ is modeled as an $\mathcal{F}_{t}$-adapted counting process, defined as
@@
The statuses of the contracts are represented by the vector $\Xi_{t}=\left(\xi_{t}^{(1)},\ldots,\xi_{t}^{(N_{t})}\right)$. Denote by $Z_{i}=\mathbb{I}_{A_{T_{i}}}l_{i}$ the payout of insurance contract $i$, where $(\mathbb{I}_{A_{T_{i}}})_{i=1,\ldots,N_{t}}$ is a sequence of Bernoulli variables, $\text{Bernoulli}(\theta_{i})$, for $i=1,\ldots,N_{t}$.
@@
$L_{0}=0,\quad L_{t}=\sum_{i=1}^{N_{t}}Z_{i}\mathbb{I}_{S_{i}\leq t,T_{i}>t},\quad\text{for }t\geq 0,$ (2)
it corresponds to the future losses given the open contracts in our portfolio at time $t$. The premiums collected to compensate for this liability are given by
@@
$\pi_{i}=(1+\eta_{S_{i}})\cdot\theta_{i}\cdot l_{i},\quad\text{for }i=1,\ldots,n.$
@@
$SCR_{t}=\text{Quantile}(L_{t}-\Pi_{t};\alpha_{\text{SCR}}),\quad\text{ and }MCR_{t}=\text{Quantile}(L_{t}-\Pi_{t};\alpha_{\text{MCR}}).$
@@
Here, $SCR$ stands for a simplified version of Solvency Capital Requirement, and $MCR$ stands for Minimum Capital Requirement, by analogy with the European directive Solvency II, under which $\alpha_{\text{SCR}}=0.995$ and $\alpha_{\text{MCR}}=0.85$.
2.2.2 Balance of the smart contract and token holdings
@@
When deployed on the Ethereum blockchain, a smart contract is assigned an Ethereum address, enabling it to receive and send ETH. The balance $(B_{t})_{t\geq 0}$ increases each time an investor contributes funds to the contract or when a policy is underwritten and a premium is paid. Conversely, the balance decreases when an investor withdraws funds or when a compensation is paid to a policyholder in the event of a covered loss. The balance decreases when an investor withdraws some funds from the contract and when a compensation is paid to an unfortunate policyholder. Let $(\tau_{i}^{+})_{i\geq 1}$ and $(\tau_{i}^{-})_{i\geq 1}$ be the times at which investors make deposits and withdrawals, respectively. Denote the associated amounts by $(x_{i}^{+})_{i\geq 1}$ for deposits and $(x_{i}^{-})_{i\geq 1}$ for withdrawals. We have
$B_{0}=0,\text{ and }B_{t}=\sum_{i\geq 1}\mathbb{I}_{\tau_{i}^{+}\leq t}x_{i}^{+}-\sum_{i\geq 1}\mathbb{I}_{\tau_{i}^{-}\leq t}x_{i}^{-}+\sum_{i\geq 1}\mathbb{I}_{S_{i}\leq t}\pi_{i}-\sum_{i\geq 1}\mathbb{I}_{T_{i}\geq t}\mathbb{I}_{A_{T_{i}}}l_{i},\text{ for }t\geq 0.$
@@
We define $(X_{t})_{t\geq 0}$ as the surplus of the smart contract, representing the funds available to underwrite new insurance policies. Unlike the balance of the smart contract, the surplus does not increase when a new contract is underwritten. Instead, the premium is added at the time of the policy’s resolution, when it is considered "earned" rather than merely "collected". Therefore we have
$X_{0}=0,\text{ and }X_{t}=\sum_{i\geq 1}\mathbb{I}_{\tau_{i}^{+}\leq t}x_{i}^{+}-\sum_{i\geq 1}\mathbb{I}_{\tau_{i}^{-}\leq t}x_{i}^{-}+\sum_{i\geq 1}\mathbb{I}_{T_{i}\leq t}\pi_{i}-\sum_{i\geq 1}\mathbb{I}_{T_{i}\geq t}\mathbb{I}_{A_{T_{i}}}l_{i}\text{ for }t\geq 0.$
@@
Note that $B_{t}\geq X_{t}$ for every $t\geq 0$. The level of $(X_{t})_{t\geq 0}$, representing the surplus of the smart contract, is compared to $SCR_{t}$ (Solvency Capital Requirement) and $MCR_{t}$ (Minimum Capital Requirement). If $X_{t}\leq SCR_{t}$, no new policies may be underwritten, and investors are prohibited from withdrawing funds. If $X_{t}\leq MCR_{t}$, the smart contract ceases its operations, and the remaining balance is redistributed—first to the policyholders and then to the investors in a way detailed in Section 3. The active insurance policy at time of termination of the smart contract are being canceled as mentioned in Remark 2.1.
@@
Whenever the smart contract receives funds from external sources, protocol tokens are issued. Let $(M_{t})_{t\geq 0}$ denote the total number of distinct Etehereum addresses that have ever transferred funds to the smart contract, whether as investments or premium payments. Let
$Y_{t}=\left(\begin{array}[]{ccc}Y_{t}^{(1)}&\ldots&Y_{t}^{(M_{t})}\end{array}\right),$
@@
be a vector representing the token holdings of the smart contract participants. The total supply of tokens is then given by $\bar{Y}_{t}=\sum_{i=1}^{M_{t}}Y_{t}^{(i)}$. The exchange rate of tokens against ETH is modeled by a stochastic process $(r_{t})_{t\geq 0}$, defined as
$r_{0}=1,\quad r_{t}=\frac{X_{t}}{\bar{Y}_{t}},\quad\text{for }t\geq 0.$
The exchange rate must be updated every time $X_{t}$ or $\bar{Y}_{t}$ changes. If, at time $t$, an amount $x$ of funds is transferred to the smart contract, a corresponding quantity of tokens $y=x/r_{t}$ is issued. Conversely, if a quantity $y$ of tokens is burned at time $t>0$, the smart contract transfers $x=y\cdot r_{t}$ back to the transaction issuer.
Section 3 provides a detailed account of the dynamics of the various stochastic processes we have just introduced.
3 A state description of the smart contract

### v2

@@ 3.2 Parametric insurance portfolio over time
The number of contracts underwritten up to time $t\geq 0$ is modeled as an $\mathcal{F}_{t}$-adapted counting process, defined as
@@
The statuses of the contracts are stored in a vector $\Xi_{t}=\left(\xi_{t}^{(1)},\ldots,\xi_{t}^{(N_{t})}\right)$. Denote by $Y_{i}=\mathbb{I}_{Q^{i}_{T_{i}}>\bar{Q}_{i}}\cdot l_{i}$ the payout of insurance contract $i$, where $\left(\mathbb{I}_{Q^{i}_{T_{i}}>\bar{Q}_{i}}\right)_{i=1,\ldots,N_{t}}$ is a sequence of Bernoulli variables with parameters
$\theta_{i}=\mathbb{P}(Q^{i}_{T_{i}}>\bar{Q}_{i}|\mathcal{F}_{S_{i}})\text{ for}i=1,\ldots,N_{t}.$
@@
$L_{0}=0,\quad L_{t}=\sum_{i=1}^{N_{t}}Y_{i}\mathbb{I}_{S_{i}\leq t,T_{i}>t},\quad\text{for }t\geq 0.$ (2)
It corresponds to the future losses given the active policies in our portfolio at time $t$. The premiums collected to compensate for this liability are given by
@@
$\pi_{i}=(1+\eta_{S_{i}})\cdot\theta_{i}\cdot l_{i},\quad\text{for }i=1,\ldots,N_{t}.$
@@
$SCR_{t}=\text{Quantile}(L_{t}-\Pi_{t};\alpha_{\text{SCR}}),\quad\text{and}\quad MCR_{t}=\text{Quantile}(L_{t}-\Pi_{t};\alpha_{\text{MCR}}).$ (3)
@@
Here, $SCR$ stands for Solvency Capital Requirement, and $MCR$ stands for Minimum Capital Requirement, by analogy with the European directive Solvency II, under which $\alpha_{\text{SCR}}=0.995$ and $\alpha_{\text{MCR}}=0.85$.
3.3 Solvency Capital Calculation
3.3.1 Model points creation to handle dependency
The main challenge associated with calculating the risk capitals (3) lies in the study of the distribution of the liability $(L_{t})_{t\geq 0}$, more specifically the likely dependency among the Bernoulli variables. We show in 3 how to gather two correlated contracts into one within the framework of 1 and 2.
Example 3.
Consider two contracts $\left(S_{i_{1}},T_{i_{1}},Q_{i_{1}},\overline{Q}_{i_{1}},l_{i_{1}}\right)$ and $\left(S_{i_{2}},T_{i_{2}},Q_{i_{2}},\overline{Q}_{i_{2}},l_{i_{2}}\right)$ associated with the same location $Q=Q_{i_{1}}=Q_{i_{2}}$ and the same event date $T=T_{i_{1}}=T_{i_{2}}$. The two contracts do not have the same threshold nor the same underwriting time. We assume without loss of generality that $\overline{Q}_{i_{1}}\leq\overline{Q}_{i_{2}}$ and $S_{i_{1}}<S_{i_{2}}$. Denote by
$\text{MP}_{t}=\begin{cases}\{i_{1}\},&\text{if }S_{i_{1}}\leq t<S_{i_{2}},\\ \{i_{1},i_{2}\},&\text{if }S_{i_{2}}\leq t<T,\\ \emptyset,&\text{otherwise,}\end{cases}$
the time-dependent set of correlated contracts. The notation "MP" stands for Model Point and corresponds to a group of contracts. These contracts are aggregated through the definition of a payout as
$Z_{t}=\begin{cases}Y_{i_{1}},&\text{if }S_{i_{1}}\leq t<S_{i_{2}},\\ Y_{i_{1}}+Y_{i_{2}},&\text{if }S_{i_{2}}\leq t<T,\\ 0,&\text{otherwise.}\end{cases}$
The probability distribution of $Z_{t}$ for $t\geq S_{i_{2}}$ is given by
$\mathbb{P}(Z_{t}=0)=1-\theta_{i_{1}},\quad\mathbb{P}(Z_{t}=l_{i_{1}})=\theta_{i_{1}}-\theta_{i_{2}},\quad\text{and}\quad\mathbb{P}(Z_{t}=l_{i_{1}}+l_{i_{2}})=\theta_{i_{2}}.$
Following up on 3, for $t\geq 0$, we partition our insurance portfolio $\left\{\left(S_{i},T_{i},Q_{i},\overline{Q}_{i},l_{i}\right)\right\}_{i=1,\ldots,N_{t}}$ into $M_{t}$ groups $\text{MP}_{1,t},\ldots,\text{MP}_{M_{t},t}$ of correlated contracts. Each class is associated with a payout $Z_{j,t}$, $j=1,\ldots,M_{t}$. The process $(M_{t})_{t\geq 0}$ counts the number of model points defined up to time $t\geq 0$. The liability $(L_{t})_{t\geq 0}$ is rewritten as
$L_{t}=\sum_{j=1}^{M_{t}}Z_{j,t}\mathbb{I}_{T_{j}>t},$
where $T_{j}$ coincides with the common event date of the contracts in $\text{MP}_{j,t}$ and
$Z_{j,t}=\sum_{i\in\text{MP}_{j,t}}\mathbb{I}_{Q^{i}_{T_{i}}>\overline{Q}_{i}}\cdot l_{i}.$
The liability is now a sum of independent, discrete, and positive random variables. The concept of model points has been introduced in the European directive Solvency II to speed up the calculations of Best Estimate Liabilities associated with life insurance portfolios when using cash-flow projection models. Grouping and aggregating insurance contracts has become a common practice among actuaries, and various methods have been documented in the actuarial science literature, see for instance the works of Goffard and Guerrault [16], Blanchet-Scalliet et al. [5], Kiermayer and Weiß [19], Gweon et al. [18], Gweon and Li [17]. Here, the creation of model points is a workaround for the dependency of the Bernoulli variables, which allows us to study the distribution of the liability in a tractable way as shown in Section 3.3.2.
3.3.2 Recursive calculation of the solvency capitals
The liability at time $t>0$ is given by
$L_{0}=0,\quad L_{t}=\sum_{j=1}^{M_{t}}Z_{j,t}\mathbb{I}_{T_{j}>t},\quad\text{for }t\geq 0,$ (4)
where $T_{j}$ denotes the common event dates of the contracts in $\text{MP}_{j,t}$. The number of active model points is denoted by
$M_{t}^{\text{active}}=\sum_{j=1}^{M_{t}}\mathbb{I}_{T_{j}>t}\mathbb{I}_{\underset{i\in\text{MP}_{j,t}}{\min}\,S_{i}\geq t}.$
Provided that $M_{t}^{\text{active}}=n$, the payouts of $\text{MP}_{1},\ldots,\text{MP}_{n}$, at time $t\geq 0$ are denoted by $Z_{1},\ldots,Z_{n}$. Consequently, the distribution of the liability $L_{t}=\sum_{k=1}^{n}Z_{k}$ is represented as the sum of independent random variables. While it is possible to compute the exact distribution using combinatorial analysis, this method becomes impractical as $n$ increases.
A more efficient approach involves using a Fast Fourier Transform (FFT) procedure. After obtaining the probability distribution of $L_{t}$, we can compute its cumulative distribution function and inverse it using a bisection algorithm. However, despite its effectiveness, FFT cannot be implemented on the blockchain due to the limitations of smart contract functions, which are unable to encode complex operations. This constraint necessitates performing solvency capital calculations off-chain, requiring intervention from the smart contract manager. Furthermore, any changes in the portfolio at a later time $s\geq t$ would require recomputing the probability distribution of the liability $L_{s}$ to determine the Solvency Capital Requirement ($\text{SCR}_{s}$) or Minimum Capital Requirement ($\text{MCR}_{s}$).
To address this, we use a normal approximation, justified by the generalized central limit theorem, which is applicable to independent random variables with finite variance. Let $\mathbb{E}\left(Z_{k}\right)=\mu_{k}$ and $\mathbb{V}\left(Z_{k}\right)=\sigma_{k}^{2}$ for $k=1,\ldots,n$ be the mean and variance of the random variables $Z_{k}$. Define $s_{n}^{2}=\sum_{k=1}^{n}\sigma_{k}^{2}$ as the sum of these variances.
Suppose the sequence of random variables $(Z_{k})_{k=1,\ldots,n}$ satisfies the Lindeberg-Feller condition:
$\lim_{n\to\infty}\frac{1}{s_{n}^{2}}\sum_{k=1}^{n}\mathbb{E}\left[(Z_{k}-\mu_{k})^{2}\mathbb{I}_{|Z_{k}-\mu_{k}|>\epsilon s_{n}}\right]=0,\quad\text{for every }\epsilon>0.$
Then, we have the following convergence in distribution:
$\frac{1}{s_{n}}\sum_{k=1}^{n}(Z_{k}-\mu_{k})\overset{\mathcal{D}}{\longrightarrow}\mathsf{Normal}(0,1),\quad\text{as }n\to\infty,$
where $\mathsf{Normal}(0,1)$ represents the standard normal distribution and $\overset{\mathcal{D}}{\longrightarrow}$ denotes convergence in distribution.
Given that the random variables $Z_{k}$ are bounded, it is straightforward to verify that the Lyapunov condition, which is stronger than the Lindeberg-Feller condition, holds. The Lyapunov condition is satisfied if there exists a $\delta>0$ such that:
$\lim_{n\to\infty}\frac{1}{s_{n}^{2+\delta}}\sum_{k=1}^{n}\mathbb{E}\left[|Z_{k}-\mu_{k}|^{2+\delta}\right]=0.$
We note that
$L_{t}-\Pi_{t}=\sum_{k=1}^{n}\left(Z_{k}-\sum_{i\in\text{MP}_{k}}(1+\eta_{S_{i}})\cdot\theta_{i}\cdot l_{i}\right)=s_{n}\left(\frac{1}{s_{n}}\sum_{i=1}^{n}(Z_{i}-\mu_{i})\right)-\sum_{k=1}^{n}\sum_{i\in\text{MP}_{k}}\eta_{S_{i}}\cdot\theta_{i}\cdot l_{i}.$
We deduce the following approximations for the SCR and MCR:
$\text{SCR}_{t}\approx s_{n}\cdot q_{\alpha_{\text{SCR}}}-\sum_{k=1}^{n}\sum_{i\in\text{MP}_{k}}\eta_{S_{i}}\cdot\theta_{i}\cdot l_{i},\quad\text{and}\quad\text{MCR}_{t}\approx s_{n}\cdot q_{\alpha_{\text{MCR}}}-\sum_{k=1}^{n}\sum_{i\in\text{MP}_{k}}\eta_{S_{i}}\cdot\theta_{i}\cdot l_{i},$ (5)
where $q_{\alpha}$ is the quantile of order $\alpha$ of the standard normal distribution. The accuracy of the approximation in (5) depends on the convergence of the Central Limit Theorem (CLT) as $n$ grows large. A common way to improve this approximation is through the Cornish-Fisher expansion, as seen in Cornish and Fisher [9], Fisher and Cornish [14]. We limit ourselves to the third and fourth-order expansions, which involve replacing the quantile $q_{\alpha}$ by:
$q_{\alpha}+\gamma_{1}\frac{q_{\alpha}^{2}-1}{6},\quad\text{and}\quad q_{\alpha}+\gamma_{1}\frac{q_{\alpha}^{2}-1}{6}+\gamma_{2}\frac{q_{\alpha}^{3}-3q_{\alpha}}{24}-\gamma_{1}^{2}\frac{2q_{\alpha}^{3}-5q_{\alpha}}{36},$ (6)
respectively, where $\gamma_{1}$ and $\gamma_{2}$ denote the skewness and excess kurtosis of $L_{t}$, respectively. Note that the approximation in (5) is referred to as the second-order Cornish-Fisher approximation. We then decide on a threshold $\bar{n}$ such that if $n>\bar{n}$, we compute the solvency capital using the approximations in (5) or (6). Otherwise, we fall back to:
$\text{SCR}_{t}=\text{MCR}_{t}=\sum_{k=1}^{N_{t}}l_{k}.$
The determination of such a threshold can be based on a brief simulation study, as illustrated in 4.
Example 4.
We sample a synthetic parametric insurance portfolio $\{(S_{i},T_{i},Q_{i},\overline{Q}_{i},l_{i})\}_{i\geq 1}$ in order to yield exactly $n=100$ model points. To achieve this, we select $100$ combinations of location and date. For each combination, we draw uniformly at random from $\{1,2,3,\ldots,10\}$ a number of parametric insurance contracts to be aggregated in a model point. The threshold is fixed to $\overline{Q}_{i}=5$ and the compensations are drawn at random as:
$l_{i}\sim\text{Unif}(\{5,10,15,20\}),\quad\text{for }i\geq 1.$
The first five rows of such a synthetic portfolio are provided in Table 2.
S T Q l station
2024-11-23 2025-01-08 5 5 MARSEILLE-MARIGNANE
2024-12-23 2025-01-08 5 5 STRASBOURG-ENTZHEIM
2024-07-06 2025-01-15 5 5 STRASBOURG-ENTZHEIM
2024-01-26 2025-01-10 5 20 STRASBOURG-ENTZHEIM
2024-01-06 2025-01-09 5 10 MARSEILLE-MARIGNANE
Table 2: Characteristics of the first five parametric insurance contracts
We simulate 100 synthetic portfolios and calculate the relative errors between the exact and approximated solvency capitals for an increasing number of model points, $n$. The order of the quantiles are set to $q_{\alpha_{\text{MCR}}}=0.85$ and $q_{\alpha_{\text{SCR}}}=0.995$. The results of these simulations are displayed in Figure 7.
(a) Relative error on MCR
(b) Relative error on SCR
Figure 7: Relative error on MCR and SCR depending on the number of models points in the porttfolio.
The normal approximation becomes acceptable for the Minimum Capital Requirement (MCR) relatively quickly, but the relative error for the Solvency Capital Requirement (SCR) converges more slowly. Overall, adjusting for skewness is beneficial for estimating both the MCR and SCR. The level of error falls below 5% when $n\geq 15$ and drops below 1% for $n\geq 30$. If we consider a 5% error level to be acceptable, then the third-order Cornish-Fisher approximation could be used as soon as the number of model points reaches 15.
Remark 3.5.
It is common practice among insurers to limit the underwriting of highly correlated risks. This can be easily implemented by preventing underwriting if the contract is aggregated within a model point that already contains too many contracts. Note that we implicitly do this in 4, as the number of correlated risks is limited to 10. Additionally, the condition could be based on the maximum possible loss associated with the model point.
The conclusions and resulting recommendations of 4 are valid for the specific case under consideration. A similar study could be conducted to adapt the threshold for the number of model points and guide the selection of an appropriate order for the Cornish-Fisher approximation for another situation.
4 Smart contract description
A smart contract is a system whose state evolves according to the actions of blockchain users. The roles of the users of our smart contract are outlined in Section 4.1. The stochastic processes used to monitor the state of the system are defined in Section 4.2. We introduce the methods that allow users to interact with the smart contract in Section 4.3.
4.1 Role distribution
The participants in our blockchain insurance scheme include the smart contract owner, the investors (also referred to as surplus providers), and the policyholders. The processes of underwriting and claim assessment are fully automated within the smart contract protocol. The role of our investors is analogous to that of the reinsurers in Cousaert et al. [10].
4.1.1 Smart contract owner
The smart contract owner is responsible for designing the insurance protocol by writing the Solidity code for the smart contract. While deploying the smart contract on the Ethereum blockchain, the owner initializes three parameters: the premium loading $\eta$, the risk levels $\alpha_{\text{SCR}}$ and $\alpha_{\text{MCR}}$ associated with the solvency capitals, as defined in Section 3, and the lower bound $m$ for the number of active model points to use the normal approximation of Section 3.3 to calculate the solvency capitals. Additionally, the contract owner acts as the claim assessor or data oracle by inputting the observed quantity at the time of the event—in this case, the precipitation height. In the blockchain ecosystem, an oracle refers to an off-chain data provider that a smart contract can query. This feature is not implemented in our current solution, it would be beneficial to do so in order to further automate our workflow.
4.1.2 Surplus providers
Surplus providers, also known as investors, contribute financially to the smart contract to constitute the solvency capital. These contributions are denominated in ETH, the native cryptocurrency of the Ethereum blockchain. In return for their investment, investors receive tokens that represent their proportional share of the surplus. Surplus providers have the option to withdraw a portion of their funds by burning these tokens in exchange for an equivalent amount of ETH, with the exchange rate defined in the Section 4.2. Withdrawals are permitted provided that the funds remaining in the smart contract do not fall below a specified threshold, known as the Solvency Capital Requirement (SCR).
4.1.3 Policyholders
Policyholders seek financial protection against adverse events through our smart contract. Upon the occurrence of such events, the smart contract pays a lump-sum compensation in ETH. To obtain this coverage, policyholders must pay a premium, also in ETH, to the smart contract.
An insurance agreement is only concluded if the Solvency Capital Requirement (SCR), after accounting for the new risk, does not exceed the funds currently locked in the smart contract. Should the settlement of a claim lead to the smart contract’s funds falling below the Minimum Capital Requirement (MCR), the system enters a state of bankruptcy. In such instances, priority is given to reimbursing policyholders with the remaining funds. Any funds that remain after all policyholder reimbursements are distributed to investors, proportional to their token holdings.
4.2 State variables
The state variables of the smart contract include the stochastic processes defined in Section 3, supplemented by additional processes defined on the filtered probability space $(\Omega,\mathcal{F},(\mathcal{F})_{t\geq 0},\mathbb{P})$.
@@ 4.2 State variables
Upon deployment on the Ethereum blockchain, a smart contract is assigned an Ethereum address, enabling it to send and receive ETH. The balance $(B_{t})_{t\geq 0}$ increases when investors contribute funds or when premiums are paid for underwritten policies. Conversely, the balance decreases upon investor withdrawals or compensation payments to policyholders. Let $(\tau_{k}^{+})_{k\geq 1}$ and $(\tau_{k}^{-})_{k\geq 1}$ denote the times of investor deposits and withdrawals, respectively, with associated amounts $(x_{k}^{+})_{k\geq 1}$ for deposits and $(x_{k}^{-})_{k\geq 1}$ for withdrawals. The balance can be expressed as:
$B_{0}=0,\quad\text{and}\quad B_{t}=\sum_{k\geq 1}\mathbb{I}_{\tau_{k}^{+}\leq t}x_{k}^{+}-\sum_{k\geq 1}\mathbb{I}_{\tau_{k}^{-}\leq t}x_{k}^{-}+\sum_{i\geq 1}\mathbb{I}_{S_{i}\leq t}\pi_{i}-\sum_{i\geq 1}\mathbb{I}_{T_{i}\leq t}\mathbb{I}_{Q_{T_{i}}>\bar{Q}_{i}}l_{i},\quad\text{for }t\geq 0.$
@@
We define $(X_{t})_{t\geq 0}$ as the surplus of the smart contract, representing the funds available to underwrite new insurance policies. Unlike the balance, the surplus does not increase immediately upon underwriting a new contract. Instead, the premium is added at the policy’s resolution, when it is considered "earned." Thus, the surplus is given by:
$X_{0}=0,\quad\text{and}\quad X_{t}=\sum_{k\geq 1}\mathbb{I}_{\tau_{k}^{+}\leq t}x_{k}^{+}-\sum_{k\geq 1}\mathbb{I}_{\tau_{k}^{-}\leq t}x_{k}^{-}+\sum_{i\geq 1}\mathbb{I}_{T_{i}\leq t}\pi_{i}-\sum_{i\geq 1}\mathbb{I}_{T_{i}\leq t}\mathbb{I}_{Q_{T_{i}}>\bar{Q}_{i}}l_{i},\quad\text{for }t\geq 0.$
@@
Note that $B_{t}\geq X_{t}$ for every $t\geq 0$. The surplus $(X_{t})_{t\geq 0}$ is compared against $\text{SCR}_{t}$ and $\text{MCR}_{t}$. If $X_{t}\leq\text{SCR}_{t}$, no new policies may be underwritten, and investors are prohibited from withdrawing funds. If $X_{t}\leq\text{MCR}_{t}$, the smart contract ceases operations, and the remaining balance is redistributed—first to policyholders and then to investors, as detailed in Section 4.3.4. Active insurance policies at the time of termination are cancelled.
@@
Let $(A_{t})_{t\geq 0}$ denote the total number of distinct Ethereum addresses that have transferred funds to the smart contract and received protocol tokens. Let
$Y_{t}=\begin{pmatrix}Y_{t}^{(1)}&\ldots&Y_{t}^{(A_{t})}\end{pmatrix},$
@@
be a vector representing the token holdings of the smart contract participants. The total supply of tokens is given by $\overline{Y}_{t}=\sum_{i=1}^{A_{t}}Y_{t}^{(i)}$. The exchange rate of tokens against ETH is modeled by the stochastic process $(r_{t})_{t\geq 0}$, defined as:
$r_{0}=1,\quad r_{t}=\frac{X_{t}}{\overline{Y}_{t}},\quad\text{for }t\geq 0.$
The exchange rate must be updated whenever $X_{t}$ or $\overline{Y}_{t}$ changes. Lastly, denote by
$\Lambda_{0}=0,\quad\Lambda_{t}=\sum_{i=1}^{N_{t}}l_{i}\mathbb{I}_{S_{i}\leq t,T_{i}>t},\quad\text{for }t\geq 0,$
the sum of all the potential compensations to be paid.
4.3 Method of the smart contract